# GAS: Generative Avatar Synthesis from a Single Image

### v1

@@ 4.4 Ablation Studies and Analyses
We conduct ablation studies evaluating variants of our proposed method. Please refer to the supplement for more ablations with quantitative and qualitative comparisons.
Generalizable geometry and appearance cues. To study the effect of both geometry (i.e., SMPL normal map) and appearance cues (i.e., human radiance field), we train a variant with the geometry cue completely removed and a variant with appearance cue removed. Our hypothesis for the generalizable human radiance field is that it provides rich and consistent appearance cues across different views and times. To validate this, we present quantitative comparisons for novel view synthesis on the THuman dataset and pose animation from a novel view on the MVHumanNet dataset. As shown in Table 3, our approach consistently improved performance across both tasks and datasets.
Metric: FVD $\downarrow$ THuman MVHumanNet
w.o. appear. cue 207.5 234.6
w. appear. cue 194.8 188.5
Table 3: Quantitative ablation study on our appearance cue. Please refer to Table 10 in the supplement for a complete quantitative evaluation on geometry and appearance cues.
@@
Figure 9: Ablation on the switcher for disentangling static view and dynamic motion synthesis. Without the switcher, undesired clothing deformations—such as dress swinging—are involved in the static novel view generation.
To demonstrate the importance of the geometry cue, we present a qualitative ablation study on internet videos (e.g., TikTok dataset). Occlusion is a common challenge in in-the-wild videos, often causing artifacts in generalizable human radiance fields due to pixel-aligned feature extraction and warping. As shown in Figure 6, without a clean geometry guidance, the radiance field can misguide the diffusion generation, leading to visible distortions.
Mixed training with 3D/multi-view captures and internet videos. We analyze the impact of our training strategy, which leverages diverse data sources, aiming to improve generalizability. As shown in Figure 7, our model demonstrates strong generalization capability for novel view synthesis in real-world scenarios. Figure 8 presents an ablation on involving internet videos for training.
Effect of switcher for disentangled synthesis. In our setup, where we jointly learn static view synthesis and dynamic motion modeling, we introduce a switcher module. To illustrate the effectiveness of this switcher, we train a model variant without the switcher and present qualitative ablation studies. Without the switcher, dynamic motion often appears in the generated consecutive novel views, as shown in Figure 9.
@@ 5 Conclusion
We present a generalizable and unified framework for avatar synthesis from a single image, addressing key challenges in appearance consistency and generalization to in-the-wild avatars. Our approach integrates a regression-based human radiance field with a video diffusion model, effectively utilizing dense conditioning to minimize discrepancies between driving signals and target representations. Additionally, our framework enables large-scale training on realistic human data from different sources, while maintaining a disentangled modeling of static novel views and dynamic motion. This synergy leads to high-fidelity view synthesis and pose animation with lifelike deformations, demonstrating strong generalization across diverse real-world scenarios.
@@ A Limitations and future works
Although our method achieves state-of-the-art performance in both novel view and pose synthesis, it is not free from limitations. (1) Due to the constraints of current off-the-shelf single-image human mesh recovery methods, we use SMPL [30] to establish geometry and appearance conditions for scalable training. However, SMPL lacks expressiveness in regions such as the face and hands, resulting in artifacts in these areas. Future works could explore scalable solutions inspired by recent efforts, such as regional supervision [51, 58]. (2) While we utilize Stable Video Diffusion (SVD) [2] to model view and pose synthesis, achieving strong consistency across both spatial and temporal dimensions, some challenges remain. Specifically, we occasionally observe degraded frame quality and difficulties in accurately generating complex clothing textures, particularly during significant clothing deformations. Addressing the former may involve exploring advanced video generative models, such as Diffusion Transformers [33, 53]. For the latter, leveraging synthetic human datasets with intricate textures presents an exciting avenue for future research.
@@ C.1 Model architecture
For the geometry cue, represented as a sequence of SMPL [30] normal maps, we use a 2D ConvNet $\bm{\epsilon}_{\text{geo}}$ to extract features $C_{\text{smpl}}\in\mathbb{R}^{T\times H_{1}\times W_{1}\times C_{1}}$. Similarly, for the appearance cue, which consists of corresponding NeRF renderings, we pass the sequence through a VAE and subsequently a 2D ConvNet $\bm{\epsilon}_{\text{appr}}$ to obtain features $C_{\text{nerf}}\in\mathbb{R}^{T\times H_{1}\times W_{1}\times C_{1}}$.
@@ D Additional results
D.1 Video results
  • In-the-wild avatar synthesis: We demonstrate novel view synthesis, pose animation and interactive 4D synthesis on in-the-wild avatars.
D.2 Runtime at inference
@@
Novel pose synthesis. We report the runtime for generating 100 consecutive novel poses from a single input image, as shown in Table 5. The additional video diffusion inference time arises due to the 6-frame overlap between consecutive video segment windows [58, 41].
@@
D.3 Ablation on geometry & appearance cues
We provide a complete quantitative ablation study on the geometry and appearance cue, as shown in Table 10. We use THuman [55] to evaluate the multi-view synthesis and MVHumanNet [50] to evaluate the novel view animation performance.
D.4 Ablation on merging novel view & pose tasks
We provide ablation analysis on merging both novel view and novel pose tasks into one model. From the static novel view synthesis perspective, the unified framework allow us to train on the abundant internet videos - which leads to the improved generalizability, as shown in Figure 8. It does not hurt the in-domain dataset performance according to Table 9.
@@ D.4 Ablation on merging novel view & pose tasks
D.5 Ablation on switcher
In addition to the qualitative ablation for the switcher presented in the main paper and video results, we also conduct quantitative ablation analysis on the switcher. With the switcher, our method effectively supports both novel view and pose synthesis, while also providing comparable or even better quantitative results, as shown in Table 8.
@@ D.5 Ablation on switcher
D.6 Novel pose synthesis on scan datasets
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
THuman MVHumanNet THuman MVHumanNet THuman MVHumanNet THuman MVHumanNet
w.o. geo. cue 26.07 28.33 0.938 0.943 0.045 0.044 227.1 210.6
w.o. appear. cue 26.38 27.66 0.938 0.941 0.041 0.043 207.5 234.6
Ours 26.77 28.74 0.943 0.945 0.041 0.040 194.8 188.5
Table 10: Quantitative ablation on geometry and appearance cues.
D.7 Robustness to input view angles
@@ E Verification of design choices
Image diffusion model is believed to be good at image-to-image translation tasks. Specifically, given a single reference image and the target appearance cue, we can train a image diffusion model to synthesize the target image. The model architecture can be a variant of Animate Anyone [16] and Champ [63], where we adopt two ReferenceNets to inject the rich information from the reference image and appearance cue respectively. The results are shown in Figure 12. To enable multi-view consistent synthesis for a single subject, we have tried adding an 1D temporal-axis attention layers [12] and only fine-tune these new added layers. We have also attempted to leverage human geometric prior to construct 3D correspondence across different views [32, 18].
Figure 12: Image-to-image novel view synthesis on MVHumanNet [50] dataset by using image diffusion model and appearance cue.
@@ E.2 Novel view camera trajectories
On our 3D scan dataset, we render a smooth camera trajectory with the same elevation and evenly distributed azimuth. We also explore the possibility of leveraging the fluctuant novel views in MVHumanNet dataset [50] to learn the multi-view consistency. However, we empirically find that the diffusion models are not able to capture the consistency even with the aid of appearance cues, causing the textures to ’flow’ across views. The reason might be two folds: 1) number of cameras is too sparse in our adopted subset of MVHumanNet dataset; 2) diffusion models cannot establish accurate correspondence for arbitrary two views with human geometric prior.

### v2

@@ 4.4 Ablation Studies and Analyses
We perform ablation studies to assess different variants of our proposed method. Additional quantitative and qualitative results can be found in the supplementary material.
Generalizable geometry and appearance cues. To study the effect of both geometry (i.e., SMPL normal map) and appearance cues (i.e., human radiance field), we train a variant with the geometry cue completely removed and a variant with appearance cue removed. We present quantitative comparisons for novel view synthesis on the THuman dataset and novel pose animation on the MVHumanNet dataset. As shown in Table 3, our approach consistently improved performance across both tasks and datasets.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
NVS NPS NVS NPS NVS NPS NVS NPS
w.o. geo. cue 26.07 28.33 0.938 0.943 0.045 0.044 227.1 210.6
w.o. appear. cue 26.38 27.66 0.938 0.941 0.041 0.043 207.5 234.6
Ours 26.77 28.74 0.943 0.945 0.041 0.040 194.8 188.5
Table 3: Quantitative ablation study on the geometry and appearance cues. NVS denotes novel view synthesis and NPS denotes novel pose synthesis. The same notation applies to Table 4.
Our hypothesis for the generalizable human NeRF is that it provides rich and consistent appearance cues across different views and times. To validate this, we present an additional qualitative study, as suggested in Figure 5.
@@
Occlusions, which are frequent in in-the-wild videos, often introduce artifacts in generalizable human NeRF due to pixel-aligned feature extraction and warping. As illustrated in Figure 6, the absence of clean geometric guidance can mislead the diffusion model, resulting in noticeable distortions.
Video diffusion model. The video diffusion model is suggested to refine the generalizable human NeRF renderings and produce shaper results with fewer artifacts. We ablate the effect of video diffusion model in Table 4 by comparing the generalizable human NeRF renderings and the diffusion refined results. Results show significant improvement across all metrics on both tasks.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
NVS NPS NVS NPS NVS NPS NVS NPS
Before diffusion 24.25 18.37 0.925 0.809 0.073 0.233 517.7 1255.8
After diffusion 28.82 19.11 0.954 0.833 0.039 0.176 191.3 362.0
Table 4: Quantitative ablation on the video diffusion model.
Mixed training with 3D/multi-view captures and internet videos. We evaluate the effectiveness of our training strategy that incorporates diverse data sources to enhance generalization. As illustrated in Figure 7, our model exhibits robust performance in novel view synthesis across real-world scenarios. Additionally, Figure 8 shows an ablation study highlighting the contribution of internet video data to improved generalization.
@@
Figure 9: Ablation on the switcher for disentangling static view and dynamic motion synthesis. Without the switcher, undesired clothing deformations, such as dress swinging, are involved in the novel view generation, corrupting the view consistency.
Effect of switcher for disentangled synthesis. In our unified framework for static view synthesis and dynamic motion modeling, we incorporate a switcher module to explicitly separate the two tasks. To assess its impact, we train a variant without the switcher. As shown in Figure 9, omitting the switcher leads to unintended motion artifacts in sequences of novel views, highlighting its importance for task disentanglement.
@@ 5 Conclusion
We propose a unified framework for avatar synthesis from a single image that tackles two major challenges: maintaining appearance consistency and generalizing to in-the-wild scenarios. Our method combines a regression-based human radiance field with a video diffusion model, leveraging dense conditioning to reduce the mismatch between driving signals and target outputs. The framework supports large-scale training on diverse, real-world human data, while cleanly separating the modeling of static novel views and dynamic motion. This combination enables high-fidelity avatar generation with realistic deformations and strong generalization across varied environments.
Acknowledgements
This work was supported by the Shanghai Artificial Intelligence Laboratory. The authors would like to thank Aviral Chharia, Jialu Gao, Jianjin Xu, and Wenbo Gou for their valuable feedback on the manuscript.
@@ A Limitations and future works
Although our method achieves state-of-the-art performance in both novel view and pose synthesis, it is not free from limitations. (1) Due to the constraints of current off-the-shelf single-image human mesh recovery methods, we use SMPL to establish geometry and appearance conditions for scalable training. However, SMPL lacks expressiveness in regions such as the face and hands, resulting in artifacts in these areas. Future works could explore scalable solutions inspired by recent efforts, such as regional supervision [55]. (2) While we utilize Stable Video Diffusion to model view and pose synthesis, achieving strong consistency across both spatial and temporal dimensions, some challenges remain. Specifically, we occasionally observe degradation in fine-grained detail quality and difficulties in accurately generating complex clothing textures, particularly during significant clothing deformations. Addressing the former may involve increasing the image resolution and exploring advanced video generative models, such as CogVideoX [58]. For the latter, leveraging synthetic human datasets with intricate textures presents an exciting avenue for future research.
@@ C.1 Model architecture
For the geometry cue, represented as a sequence of SMPL normal maps, we use a 2D ConvNet $\bm{\epsilon}_{\text{geo}}$ to extract features $C_{\text{smpl}}\in\mathbb{R}^{T\times H_{1}\times W_{1}\times C_{1}}$. Similarly, for the appearance cue, which consists of corresponding NeRF renderings, we pass the sequence through a VAE and subsequently a 2D ConvNet $\bm{\epsilon}_{\text{appr}}$ to obtain features $C_{\text{nerf}}\in\mathbb{R}^{T\times H_{1}\times W_{1}\times C_{1}}$.
@@ D Additional results
D.1 Runtime at inference
@@
Novel pose synthesis. We report the runtime for generating 100 consecutive novel poses from a single input image, as shown in Table 6. The additional video diffusion inference time arises due to the 6-frame overlap between consecutive video segment windows.
@@
D.2 Ablation on merging novel view & pose tasks
We provide quantitative ablation analysis on merging both novel view and novel pose tasks into one model. From the static novel view synthesis perspective, the unified framework allow us to train on the abundant internet videos - which leads to the improved generalization, as shown in the main paper. It does not hurt the in-domain dataset performance according to Table 10.
@@ D.2 Ablation on merging novel view & pose tasks
D.3 Ablation on switcher
In addition to the qualitative ablation for the switcher presented in the main paper, we also conduct quantitative ablation analysis on the switcher. With the switcher, our method effectively supports both novel view and pose synthesis, while also providing comparable or even better quantitative results, as shown in Table 9.
@@ D.3 Ablation on switcher
D.4 Novel pose results on scan datasets
@@
D.5 Robustness to input view angles
@@
D.6 Ablation on video diffusion model
Fig. 12 (left) presents quantitative ablations on both novel view and pose synthesis tasks, showing consistent improvements over the direct output of the generalizable NeRF results, denoted as before diff. Fig. 12 (right) shows qualitative results on MVHumanNet dataset: NeRF helps the diffusion model preserve identity on nearby views, while the diffusion model refines distant views, producing sharper results with fewer artifacts.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
NVS NPS NVS NPS NVS NPS NVS NPS
Before diff. 24.25 18.37 0.925 0.809 0.073 0.233 517.7 1255.8
After diff. 28.82 19.11 0.954 0.833 0.039 0.176 191.3 362.0
Figure 12: Quantitative and qualitative results before/after diff.
D.7 Video results
  • In-the-wild avatar synthesis: We demonstrate novel view synthesis, pose animation and interactive 4D video synthesis on in-the-wild avatars.
@@ E Verification of design choices
Image diffusion model is believed to be good at image-to-image translation tasks. Specifically, given a single reference image and the target appearance cue, we can train a image diffusion model to synthesize the target image. The model architecture can be a variant of Animate Anyone [16] and Champ [67], where we adopt two ReferenceNets to inject the rich information from the reference image and appearance cue respectively. The results are shown in Figure 13. To enable multi-view consistent synthesis for a single subject, we have tried adding an 1D temporal-axis attention layers [12] and only fine-tune these new added layers. We have also attempted to leverage human geometric prior to construct 3D correspondence across different views [19].
Figure 13: Image-to-image novel view synthesis on MVHumanNet dataset by using image diffusion model and appearance cue.
@@ E.2 Novel view camera trajectories
On our 3D scan dataset, we render a smooth camera trajectory with the same elevation and evenly distributed azimuth. We also explore the possibility of leveraging the fluctuant novel views in MVHumanNet dataset to learn the multi-view consistency. However, we empirically find that the diffusion models are not able to capture the consistency even with the aid of appearance cues, causing the textures to flow across views.